find area if length of a rectangle is (x+1) and its width is 5cm less than its length
Answer by Edwin McCravy(20056)

Instead of doing your problem for you, I'll do one exactly
like it so you can use it as a model to do yours exactly
the same way.  I'll do this one instead:

find area if length of a rectangle is (x+3) and its width is 9cm less than its
length

length = (x+3)
width = (x+3)-9 = x+3-9 = x-6
area = length∙width = (x+3)(x-6) = x²-6x+3x-18 = x²-3x-18

Now do yours the exact same way, step by step.  Only the numbers
are different.

Edwin